# Ada-Tracker: Soft Tissue Tracking via Inter-Frame and Adaptive-Template Matching

### v1

@@ IV-A Datasets
Method
$Rob_{2D}$$\uparrow$
$Acc_{2D}$$\uparrow$
$Err_{2D}$$\downarrow$
$Rob_{3D}$$\uparrow$
$Err_{3D}$$\downarrow$
$EAO$$\uparrow$
CSRT
0.872
0.769
5.7$\pm$2.6
0.894
3.3$\pm$3.8
0.563
Jmees
0.868
0.818 5.3$\pm$2.4
0.878
2.7$\pm$1.9
0.583
KIT
0.465
0.747
12.5$\pm$8.0
0.810
11.9$\pm$9.5
0.223
TransT
0.701
0.529
16.3$\pm$5.1
0.861
17.3$\pm$26.7
0.274
SRV
0.476
0.681
15.4$\pm$7.5
0.710
8.5$\pm$13.7
0.293
MEDCVR
0.702
0.509
7.9$\pm$5.8
0.832
9.2$\pm$39.7
0.302
ETRI
0.802
0.693
12.1$\pm$7.4
0.909
5.7$\pm$5.2
0.405
RIWOlink
0.807
0.737
8.0$\pm$4.5
0.894
5.8$\pm$9.2
0.433
ICVS-2Ai
0.872
0.816
6.7$\pm$3.5
0.901
2.6$\pm$2.3
0.573
@@ IV-C Comparison with State-of-the-art
$Rob_{2D}$$\uparrow$
$Acc_{2D}$$\uparrow$
$Err_{2D}$$\downarrow$
$Rob_{2D}$$\uparrow$
$Acc_{2D}$$\uparrow$
$Err_{2D}$$\downarrow$
$Rob_{2D}$$\uparrow$
$Acc_{2D}$$\uparrow$
$Err_{2D}$$\downarrow$
$Rob_{2D}$$\uparrow$
$Acc_{2D}$$\uparrow$
$Err_{2D}$$\downarrow$
$Rob_{2D}$$\uparrow$
$Acc_{2D}$$\uparrow$
$Err_{2D}$$\downarrow$
CSRT
0.92 0.72 7±3 0.87 0.77 8±3 0.87 0.81 5±2 1.0 0.85 3±1 0.70 0.68 6±3
Jmees
0.92 0.76 7±3 0.89 0.77 7±3 0.88 0.88 5±2 0.94 0.90 3±1 0.70 0.75 5±3
KIT
0.26 0.61 19±13 0.17 0.61 23±17 0.81 0.84 9±7 0.84 0.71 11±6 0.16 0.58 20±10
TransT
0.72 0.46 18±4 0.64 0.42 24±10 0.82 0.61 16±4 0.79 0.61 11±7 0.47 0.48 11±4
SRV
0.39 0.49 27±13 0.19 0.55 16±11 0.71 0.80 11±6 0.98 0.69 13±5 0.07 0.56 22±12
MEDCVR
0.74 0.54 7±5 0.65 0.48 16±12 0.82 0.68 4±4 0.77 0.61 10±7 0.46 0.57 6±4
ETRI
0.88 0.56 21±13 0.82 0.69 14±10 0.85 0.82 7±3 0.94 0.73 7±4 0.49 0.60 10±6
RIWOlink
0.89 0.66 9±5 0.71 0.78 9±6 0.85 0.86 3±1 0.99 0.69 10±5 0.55 0.59 9±6
ICVS-2Ai
0.94 0.76 9±4 0.84 0.79 9±6 0.91 0.88 5±3 1.0 0.90 3±1 0.65 0.72 8±5
@@
$Rob_{3D}$$\uparrow$
$Err_{3D}$$\downarrow$
$Rob_{3D}$$\uparrow$
$Err_{3D}$$\downarrow$
$Rob_{3D}$$\uparrow$
$Err_{3D}$$\downarrow$
$Rob_{3D}$$\uparrow$
$Err_{3D}$$\downarrow$
$Rob_{3D}$$\uparrow$
$Err_{3D}$$\downarrow$
CSRT
0.91 5±4 0.83 2±1 0.93 1±1 1.0 1±1 0.78 9±16
Jmees
1.0 2±1 0.85 2±1 0.93 1±1 0.95 1±1 0.72 5±4
KIT
0.68 20±16 0.70 23±14 0.97 2±2 1.0 6±5 0.68 20±19
TransT
0.91 23±8 0.82 24±53 0.92 6±14 0.92 10±56 0.70 31±26
SRV
0.66 16±13 0.37 16±12 0.96 3±3 1.0 2±1 0.46 18±66
MEDCVR
0.90 8±22 0.75 20±125 0.88 2±9 0.92 10±62 0.67 14±24
ETRI
0.97 6±4 0.90 7±6 0.94 2±3 1.0 3±3 0.71 13±13
RIWOlink
0.96 7±4 0.81 6±5 0.93 1±1 0.99 5±34 0.72 11±7
ICVS-2Ai
0.98 4±2 0.84 1±1 0.94 1±2 1.0 1±1 0.71 6±6
Ours
0.99 2±2 0.84 2±1 0.96 1±1 1.0 1±1 0.76 4±3
@@
I
T
A
$Rob_{2D}$$\uparrow$
$Acc_{2D}$$\uparrow$
$Err_{2D}$$\downarrow$
$Rob_{3D}$$\uparrow$
$Err_{3D}$$\downarrow$
$EAO$$\uparrow$
0.323
0.512
14.2$\pm$5.4
0.587
16.9$\pm$13.5
0.088
0.541
0.627
8.9$\pm$7.3
0.620
9.8$\pm$19.7
0.164
0.643
0.639
7.2$\pm$6.5
0.693
7.9$\pm$17.4
0.279

### v2

@@ IV-A Datasets
Method $Rob_{2D}$$\uparrow$ $Acc_{2D}$$\uparrow$ $Err_{2D}$$\downarrow$ $Rob_{3D}$$\uparrow$ $Err_{3D}$$\downarrow$ $EAO$$\uparrow$
CSRT 0.872 0.769 5.7$\pm$2.6 0.894 3.3$\pm$3.8 0.563
Jmees 0.868 0.818 5.3$\pm$2.4 0.878 2.7$\pm$1.9 0.583
KIT 0.465 0.747 12.5$\pm$8.0 0.810 11.9$\pm$9.5 0.223
TransT 0.701 0.529 16.3$\pm$5.1 0.861 17.3$\pm$26.7 0.274
SRV 0.476 0.681 15.4$\pm$7.5 0.710 8.5$\pm$13.7 0.293
MEDCVR 0.702 0.509 7.9$\pm$5.8 0.832 9.2$\pm$39.7 0.302
ETRI 0.802 0.693 12.1$\pm$7.4 0.909 5.7$\pm$5.2 0.405
RIWOlink 0.807 0.737 8.0$\pm$4.5 0.894 5.8$\pm$9.2 0.433
ICVS-2Ai 0.872 0.816 6.7$\pm$3.5 0.901 2.6$\pm$2.3 0.573
@@ IV-C Comparison with State-of-the-art
$Rob_{2D}$$\uparrow$ $Acc_{2D}$$\uparrow$ $Err_{2D}$$\downarrow$ $Rob_{2D}$$\uparrow$ $Acc_{2D}$$\uparrow$ $Err_{2D}$$\downarrow$ $Rob_{2D}$$\uparrow$ $Acc_{2D}$$\uparrow$ $Err_{2D}$$\downarrow$ $Rob_{2D}$$\uparrow$ $Acc_{2D}$$\uparrow$ $Err_{2D}$$\downarrow$ $Rob_{2D}$$\uparrow$ $Acc_{2D}$$\uparrow$ $Err_{2D}$$\downarrow$
CSRT 0.92 0.72 7±3 0.87 0.77 8±3 0.87 0.81 5±2 1.0 0.85 3±1 0.70 0.68 6±3
Jmees 0.92 0.76 7±3 0.89 0.77 7±3 0.88 0.88 5±2 0.94 0.90 3±1 0.70 0.75 5±3
KIT 0.26 0.61 19±13 0.17 0.61 23±17 0.81 0.84 9±7 0.84 0.71 11±6 0.16 0.58 20±10
TransT 0.72 0.46 18±4 0.64 0.42 24±10 0.82 0.61 16±4 0.79 0.61 11±7 0.47 0.48 11±4
SRV 0.39 0.49 27±13 0.19 0.55 16±11 0.71 0.80 11±6 0.98 0.69 13±5 0.07 0.56 22±12
MEDCVR 0.74 0.54 7±5 0.65 0.48 16±12 0.82 0.68 4±4 0.77 0.61 10±7 0.46 0.57 6±4
ETRI 0.88 0.56 21±13 0.82 0.69 14±10 0.85 0.82 7±3 0.94 0.73 7±4 0.49 0.60 10±6
RIWOlink 0.89 0.66 9±5 0.71 0.78 9±6 0.85 0.86 3±1 0.99 0.69 10±5 0.55 0.59 9±6
ICVS-2Ai 0.94 0.76 9±4 0.84 0.79 9±6 0.91 0.88 5±3 1.0 0.90 3±1 0.65 0.72 8±5
@@
$Rob_{3D}$$\uparrow$ $Err_{3D}$$\downarrow$ $Rob_{3D}$$\uparrow$ $Err_{3D}$$\downarrow$ $Rob_{3D}$$\uparrow$ $Err_{3D}$$\downarrow$ $Rob_{3D}$$\uparrow$ $Err_{3D}$$\downarrow$ $Rob_{3D}$$\uparrow$ $Err_{3D}$$\downarrow$
CSRT 0.91 5±4 0.83 2±1 0.93 1±1 1.0 1±1 0.78 9±16
Jmees 1.0 2±1 0.85 2±1 0.93 1±1 0.95 1±1 0.72 5±4
KIT 0.68 20±16 0.70 23±14 0.97 2±2 1.0 6±5 0.68 20±19
TransT 0.91 23±8 0.82 24±53 0.92 6±14 0.92 10±56 0.70 31±26
SRV 0.66 16±13 0.37 16±12 0.96 3±3 1.0 2±1 0.46 18±66
MEDCVR 0.90 8±22 0.75 20±125 0.88 2±9 0.92 10±62 0.67 14±24
ETRI 0.97 6±4 0.90 7±6 0.94 2±3 1.0 3±3 0.71 13±13
RIWOlink 0.96 7±4 0.81 6±5 0.93 1±1 0.99 5±34 0.72 11±7
ICVS-2Ai 0.98 4±2 0.84 1±1 0.94 1±2 1.0 1±1 0.71 6±6
Ours 0.99 2±2 0.84 2±1 0.96 1±1 1.0 1±1 0.76 4±3
@@
I T A $Rob_{2D}$$\uparrow$ $Acc_{2D}$$\uparrow$ $Err_{2D}$$\downarrow$ $Rob_{3D}$$\uparrow$ $Err_{3D}$$\downarrow$ $EAO$$\uparrow$
0.323 0.512 14.2$\pm$5.4 0.587 16.9$\pm$13.5 0.088
0.541 0.627 8.9$\pm$7.3 0.620 9.8$\pm$19.7 0.164
0.643 0.639 7.2$\pm$6.5 0.693 7.9$\pm$17.4 0.279